Session: February 1996 (session 1)
Frequency: 18 cm
Obs time (UT): 41 EU 06:00(10/02)-18:00(10/02)
EU Mk3 Tapes/St.: 3
To give feedback on this experiment, click here
Onsala (John Conway at Mon Feb 19 08:01:04 EST 1996) Minor failures
Telescope stowed between 0946UT and 1150UT to try to solve problems with driving telescope to East. Partially succesful afterwards could move at 3deg/minute in East direction (20deg/min in West direction) sufficient to keep up with schedule. Phase cal signal switched off after 0950UT, apparently producing high signal strengths in some channels and increasing system temperature.
Simeiz (Alexandre Volvatch at Wed Feb 21 08:22:58 EST 1996) Minor failures
8 passes (the end of expt) was omited because source was behind
the hill (EL=2d, 0d, -2d).
Jodrell (Dave Henstock at Thu Feb 22 06:13:26 EST 1996) Minor failures
Very bad RFI during experiment - VCs 1,3,9,10 affected most of time.
Due to RFI, Tsys measurements only crudely calibrated. Specific problems:
Day 41 09:00-16:00 RFI is broad band, appears to be in almost all VCs
16:40-17:08 and 17:21-17:33 Field system crashed resulting in no data
recorded between these times.
Tsys is in Jy, need to multiply measured Tsys by 1.23 for better calibration.
Cambridge (Dave Henstock at Thu Feb 22 06:17:31 EST 1996) Success
Apparently no problems. Note there are no Tsys measurements for Cambridge.
Tsys is 212 Jy (from EVN status).
Medicina (Venturi Tiziana at Thu Feb 22 09:43:21 EST 1996) Success
Apparently no problems
Noto (Corrado Trigilio at Sat Feb 24 06:04:09 EST 1996) Success
No problems reported. Tsys values in logfile need to be corrected; I will provvide corrected values. No phase cal.
Westerbork (foley at Mon Feb 26 11:40:19 EST 1996) Minor failures
Some shadowing at 06:33 to 07:30 UT
lost ~ 1min at 17:32 UT (system restart)
Calibration (Fredrik T Rantakyro at Thu Sep 26 10:09:32 EDT 1996) Minor failures
An ANTAB formatted file for the experiment can be found on VLBEER